OpenStudy (anonymous):

8 = 2x + a, becomes... 4 = x + a or 4 = x + a/2

OpenStudy (mattfeury):

you have to divide each side by 2. 8 / 2 = 4 (2x + a) / 2 = 2x/2 + a/2 4 = x + a/2

OpenStudy (anonymous):

ah ok, for some reason i thought that the 2 wouldn't divide every single term separately... thank you :D

OpenStudy (mattfeury):

no problem :). always remember, when you divide an equation by something, every "term" in that equation gets divided.

OpenStudy (anonymous):

writing that down verbatim